The uniform bar weighs 45 N and is subjected to the forces as shown. Find the magnitude, location and direction of the force needed to keep the bar in equilibrium


1
Expert's answer
2022-03-01T15:17:47-0500

"\Delta F_x=60+70+45-50-40=85~N,"

"\Delta F_y=69~N,"

"F=\sqrt{\Delta F_x^2+\Delta F_y^2}=110~N,"

"\alpha=\arctan\frac{\Delta F_y}{\Delta F_x}=39°."